I have two gallons of grape drink that is one-third water, and I mix it with three gallons of lemonade that is three-fourths water. After mixing, I spill one gallon of the mixture. How much water is there in the remaining four gallons of the mixture?
Answer by stanbon(75887)

I have two gallons of grape drink that is one-third water, and I mix it with three gallons of lemonade that is three-fourths water. After mixing, I spill one gallon of the mixture. How much water is there in the remaining four gallons of the mixture?
------------
Equation:
water + water = water
(1/3)2 + (3/4)3 = r(5) where r is the fraction of water in the mixture
------------------
(2/3) + (9/4) = 5r
35/12 = 5r
r = 7/12
------------------------
7/12 of the 5 gallon mixture is water
----------------
You spill 1 gallon of the mixture; you are left with
(7/12)4 = (7/3)gallon of water in the 4 gallon mixture.
====================================================
Cheers,
Stan H.